OpenStudy (anonymous):

Solve the proportion 2 times x all over 9 times x minus 8 equals 2 over x.. What is the value of x? x = 1 and x = 9 x = −1 and x = −9 x = 1 and x = 8 x = −1 and x = 8

OpenStudy (texaschic101):

2x/(9x - 8) = 2/x cross multiply (2x)(x) = 2(9x - 8) 2x^2 = 18x - 16 2x^2 - 18x + 16 = 0 2(x - 8)(x - 1) = 0 x - 8 = 0 x = 8 x - 1 = 0 x = 1 I am not sure if I did this correctly...I am not good at quadratic equations...you will need to get a second opinion.

OpenStudy (anonymous):

well thank you texas chi

OpenStudy (texaschic101):

like I said...you might want to get a second opinion

OpenStudy (anonymous):

ok.

jimthompson5910 (jim_thompson5910):

x = 8 or x = 1 are the two solutions, so you did it correctly texaschic101

OpenStudy (texaschic101):

I am glad....I kinda guessed....I am not good at quadratic equations

jimthompson5910 (jim_thompson5910):

you did good, just have to trust yourself more

OpenStudy (anonymous):

(sharun) hey everyone watch out for this user i sent a report and blocked him but he is really uncomfortably flirtatious.

OpenStudy (texaschic101):

thanks for the warning :)

jimthompson5910 (jim_thompson5910):

if he gets out of control, just report him and the mods will know what to do with him

jimthompson5910 (jim_thompson5910):

sadly a lot of people think this is a dating website (not sure why)